# How to choose efficiently the size of the Bethe-Salpeter Equation Hamiltonian for accurate exciton calculations on supercells

### v2

@@ II Theory
    Perform a converged GW/BSE calculation for the PC with a fine $k$-grid $N_{1}N_{k,1}\times N_{2}N_{k,2}\times N_{3}N_{k,3}$. Save the QP energy levels on this fine grid, the eigenvectors file with the $A_{cv\boldsymbol{k}}$ coefficients, and the BSE Hamiltonian matrix elements. All those files can be saved in the absorption step of the BerkeleyGW code [27], version 4.0. If the BSE Hamiltonian file is not available, as for earlier versions or other codes, the BSE Hamiltonian can be reconstructed using the exciton energies and coefficients through $\langle c^{\prime}v^{\prime}\boldsymbol{k}^{\prime}|H^{\rm{BSE}}|cv\boldsymbol{k}\rangle=\sum_{A}\Omega_{A}A^{*}_{c^{\prime}v^{\prime}\mathbf{k^{\prime}}}A_{cv\boldsymbol{k}}$.
@@
    Use the exciton coefficients $A_{cv\mathbf{k}}$ and QP energy levels from the PC to estimate the BSE size for the desired precision of the binding energy in SC calculations. Equation 2 provides an estimate of the necessary energy window, and Eqs. 4 and 5 can be used to predict the error in the exciton energy (Fig. 4). In Fig. 2, we show how the partial exciton energy (Eq. 4) depends on the number of conduction and valence bands. We provide a Jupyter notebook and example data for this analysis [54] that is compatible with BerkeleyGW [27].
Figure 2: First exciton energy evolution as function of the number of valence ($N_{v}$) and conduction ($N_{c}$) bands used to build the BSE Hamiltonian. Exciton energy is calculated using Eq. 4.
@@ III Computational Details
The PC BSE Hamiltonian used for the calculation results is built with 5 valence and 10 conduction bands. Only 3 valence and 1 conduction bands are necessary to converge the first optical absorption peaks, and we present later timing comparisons from this minimal set of bands, but we include more bands in other calculations to allow a more detailed exploration of a wider range in the spectrum. Optical absorption plots reported in this paper were calculated with light polarization parallel to the direction $-\hat{x}+\hat{y}$, with the momentum operator [27, 25] (which does not include non-local effects) for simplicity, as our objective in this work is only to study the convergence.
@@ IV.2 How to choose the BSE size for SC calculations
Next, we estimate SC results from PC results focusing on the first and second absorption peaks (Fig. 5(a)). In Fig. 5(b) and (c), we plot the expected binding energy variation with respect to PC results for the first and second absorption peaks, respectively. The expected binding energy was given by the expected exciton energy from Eq. 4, subtracted from the QP energy gap. In our code (provided at [54]), we count how many valence and conduction states $N_{c}$, $N_{v}$ are present in a given energy window $[E^{\rm{min}},E^{\rm{max}}]$. Then we estimate the value of $\Omega^{\rm part}_{A}$ for the BSE size $(N_{c}N_{v}N_{k}^{2})^{2}$. The expected binding energy increases roughly monotonically with the BSE size until it reaches the PC results when the BSE size is equal to the size of a BSE Hamiltonian obtained by zone folding. The reference binding energies for the PC can be found in Table 2.
To test the quality of the predictions made by PC calculations, we performed SC calculations with different BSE Hamiltonian sizes, and the agreement between the two data sets is excellent (Fig. 5 (b) and (c)). This shows the efficiency of our analysis in predicting results on BSE SC calculations based on the results of PC. In our final result for SC calculations, the BSE is composed of 137 valence bands and 31 conduction bands with 1 $k$-point (just $\Gamma$), which means that $(1\times 137\times 31)^{2}\sim 18\times 10^{6}$ matrix elements are necessary. By zone-folding 3 valence and 1 conduction bands of a $k$-grid with 64 $k$-points of the PC into the SC Brillouin zone, the SC calculation would need 3$\times$64 valence, 64 conduction bands and 1 $k$-point, which means that $(3\times 64\times 64\times 1)^{2}\sim 151\times 10^{6}$ matrix elements would be necessary. Therefore, our approach requires only 12% of the matrix elements one would expect to need from only using zone folding. The error is 150 meV for the first exciton, which is 1.2% of the first excitation energy value. In Table 1, we show some time estimation comparisons, showing how our approach can save around a factor of 10 in the time for SC BSE calculations, given the reduction in the number of matrix elements by 12%. Next, we show the optical absorption spectra for PC and SC cases in Fig. 5 (a). We observed that for energies up to 17 eV, BSE absorption peaks for SC are blueshifted by 0.15 eV in relation to PC results (Table 2), while no absorption peak is present for energies above 17 eV in the SC BSE spectrum with the limited number of bands used.
@@ IV.3 Relation of excitonic properties with BSE size
V Conclusions
We presented in this paper an effective scheme to set the BSE Hamiltonian size for SC calculations based on PC results without the need for new convergence studies, which we demonstrated with detailed analysis of the case of rocksalt LiF. The scheme is applicable to exact SCs or cases of slightly perturbed atomic positions, in which case analysis of finite-momentum excitons in the PC is likely also needed, to converge the unperturbed JDOS as well as the absorption spectrum. We introduced the Weighted Average Single-Particle Energies (WASPEs) for conduction and valence of each exciton, to indicate the scale of energies that need to be included in the SC BSE Hamiltonian, and we further analyzed the energy window of states participating significantly, which may extend for another 1-2 eV, to determine the states needed. We found a reduction of an order of magnitude in the number of matrix elements needed, leading to an order of magnitude reduction in computational expense for BSE. Our analysis shows that the kernel matrix calculation is the key concern since its expense rises more rapidly with supercell size than the absorption calculation in which the matrix is diagonalized. Our approach for determining the number of states needed in BSE also allows one to choose for which energy levels GW calculations are necessary, avoiding unnecessary calculations and making SC GW/BSE calculations more accessible. We provide a Jupyter notebook and example data [54] with the analysis from PC results that can be used for materials in general. This method may be applied to study the optical properties of SC models such as surfaces, defects, and polarons in diverse materials, speeding up ab initio results and decreasing RAM demands of exciton calculations substantially, and enabling studies of more complex quantum effects in materials.

### v3

@@ II.3 Workflow
    Perform a converged GW/BSE calculation for the PC with a fine $k$-grid $N_{1}N_{k,1}\times N_{2}N_{k,2}\times N_{3}N_{k,3}$. Save the QP energy levels on this fine grid, the eigenvectors file with the $A_{cv\boldsymbol{k}}$ coefficients, and the BSE Hamiltonian matrix elements. All those files can be saved in the absorption step of the BerkeleyGW code [31], version 4.0 (using the flags dump_bse and write_eigenvectors of the Absorption step). If the BSE Hamiltonian file is not available, as for earlier versions or other codes, the BSE Hamiltonian can be reconstructed using the exciton energies and coefficients through $\langle c^{\prime}v^{\prime}\boldsymbol{k}^{\prime}|H^{\rm{BSE}}|cv\boldsymbol{k}\rangle=\sum_{A}\Omega_{A}A^{*}_{c^{\prime}v^{\prime}\mathbf{k^{\prime}}}A_{cv\boldsymbol{k}}$.
@@
    Use the exciton coefficients $A_{cv\mathbf{k}}$ and QP energy levels from the PC to estimate the BSE size for the desired precision of the binding energy in SC calculations. Equation 2 provides an estimate of the necessary energy window, and Eqs. 4 and 5 can be used to predict the error in the exciton energy (Fig. 4). In Fig. 2, we show how the partial exciton energy (Eq. 4) depends on the number of conduction and valence bands. Alternatively, one can plot how much of the exciton weight is preserved inside the chosen energy window, because this quantity has a quadratic relation with the partial exciton energy (Fig. 2(b)). We provide a Jupyter notebook and example data for this analysis [65] that is compatible with BerkeleyGW [31]. The Jupyter notebook also has the code necessary to recover the BSE Hamiltonian from the BSE eigenvectors using the equation in the end of step 2.
Figure 2: (a) First exciton energy evolution, in comparison to the full PC BSE calculation $\Omega^{\rm{PC}}_{1}$, as function of the number of valence ($N_{v}$) and conduction ($N_{c}$) bands used to build the BSE Hamiltonian. Exciton energy is calculated using Eq. 4. (b) Correlation of $\Omega^{\rm{part}}-\Omega^{\rm{PC}}_{1}$ with the partial sum of the exciton coefficients, following a quadratic relation.
The present workflow can be applied to any exciton of interest (specified by its index). Our Jupyter notebook runs in just a few minutes in a personal computer, allowing one to explore the convergence of several excitons. In our case of LiF, both excitons are localized at $\Gamma$ and their convergence is pretty similar (Fig. 5). By contrast, in monolayer MoS2, the A exciton is localized at K and the C exciton is composed of transitions at K and around $\Gamma$ [55]. In this case, one needs to make the convergence study for each exciton separately and use the larger of the number of bands in order to converge both simultaneously. In general, higher excitons may spread more in energy and in $k$ space, which would require more bands to be converged. These calculations are done by selecting the same number of valence (conduction) bands per k-point, which is what is done in BerkeleyGW as part of the memory layout to enable efficient parallel matrix-vector operations [31]. Alternate schemes in which arbitrary pairs of states are selected could be used to reduce the basis further but would be inconsistent with this layout for computational performance. Trying to use folding back to a PC BZ in order to exclude transitions between different PC $\mathbf{k}$ points is unlikely to be helpful because cases of interest for a SC calculation are the ones where there is a breaking of symmetry and the mixing of states from different $\mathbf{k}$-points may be important for the optical properties of interest, e.g. brightening of dark excitons, so such an approach could save some computational time but some of the physics may be lost as well.
@@ III Computational Details
The PC BSE Hamiltonian used for the calculation results is built with 5 valence and 10 conduction bands. Only 3 valence and 1 conduction bands are necessary to converge the first optical absorption peaks, and we present later timing comparisons from this minimal set of bands, but we include more bands in other calculations to allow a more detailed exploration of a wider range in the spectrum. Optical absorption plots reported in this paper were calculated with light polarization parallel to the direction $\hat{x}+\hat{y}$, with the momentum operator [31, 25] (which does not include non-local effects) for simplicity, as our objective in this work is only to study the convergence.
@@ IV.2 How to choose the BSE size for SC calculations
Next, we estimate SC results from PC results focusing on the first and second absorption peaks (Fig. 5(a)). In Fig. 5(b) and (c), we plot the expected binding energy variation with respect to PC results for the first and second absorption peaks, respectively. The expected binding energy was given by the expected exciton energy from Eq. 4, subtracted from the QP energy gap. In our code (provided at [65]), we count how many valence and conduction states $N_{c}$, $N_{v}$ are present in a given energy window $E^{\rm{min}},E^{\rm{max}}]$. Then we estimate the value of $\Omega^{\rm part}_{A}$ for the BSE size $(N_{c}N_{v}N_{k}^{2})^{2}$. The expected binding energy increases roughly monotonically with the BSE size until it reaches the PC results when the BSE size is equal to the size of a BSE Hamiltonian obtained by zone folding. The reference binding energies for the PC can be found in Table 2. Beyond the lowest energy excitons, the energy window to ensure convergence increases, so more bands are necessary to build the BSE Hamiltonian as shown in Fig. 4.
[…]
@@ IV.3 Relation of excitonic properties with BSE size
V Vk-like defect in LiF
As a proof of concept we performed a GW/BSE calculation for a Vk-center-like defect in LiF [77], and our absorption spectra results are shown in Fig. 8. Vk centers in alkali halides are defects where the halide atoms get closer to each other, similar to what is shown in Fig. 8. To represent well the Vk center one would need to include the spin degree of freedom and relax the excited state [10] for the defect exciton, which we did not do as it is out of the scope of this work. This result illustrates how defect calculations can be performed using our scheme. It uses the same parameters of our SC BSE calculations reported in Table 2, notably the number of valence and conduction bands used in the BSE, which was provided by our method. Two fluorine atoms are moved towards each other in the [110] direction and the distance between them is 1.4 Å. This is a somewhat arbitrary displacement to break the symmetry which can be used for further excited-state relaxations[10]. In our BSE results we can observe that the two lowest peaks of the case without defect are still present, and the second lowest peak is redshifted. In addition to that, new peaks between 6 and 12 eV appear, which are transitions including mid gap defect states.
Figure 8: Optical absorption for rocksalt (RS) LiF and a Vk-like defect, both calculated in a $\times4$$\times$4 SC. Inset: Structure of the defect in which 2 fluorine atoms are moved towards each other in the [110] direction.
VI Conclusions
We presented in this paper an effective scheme to set the BSE Hamiltonian size for SC calculations based on PC results without the need for new convergence studies, which we demonstrated with detailed analysis of the case of rocksalt LiF. The scheme is applicable to exact SCs or cases of slightly perturbed atomic positions, in which case analysis of finite-momentum excitons in the PC is likely also needed, to converge the unperturbed JDOS as well as the absorption spectrum. We introduced the Weighted Average Single-Particle Energies (WASPEs) for conduction and valence of each exciton, to indicate the scale of energies that need to be included in the SC BSE Hamiltonian, and we further analyzed the energy window of states participating significantly, which may extend for another 1-2 eV, to determine the states needed. We found a reduction of an order of magnitude in the number of matrix elements needed, leading to an order of magnitude reduction in computational expense for BSE. Our analysis shows that the kernel matrix calculation is the key concern since its expense rises more rapidly with supercell size than the absorption calculation in which the matrix is diagonalized. Our approach for determining the number of states needed in BSE also allows one to choose for which energy levels GW calculations are necessary, avoiding unnecessary calculations and making SC GW/BSE calculations more accessible. Then we showed, as a proof of concept, BSE calculations on a Vk-like defect in LiF, demonstrating the applicability of our method. We provide a Jupyter notebook and example data [65] with the analysis from PC results that can be used for materials in general. This method may be applied to study the optical properties of SC models such as surfaces, defects, and polarons in diverse materials, speeding up ab initio results and decreasing RAM demands of exciton calculations substantially, and enabling studies of more complex quantum effects in materials, and can be used within any code that builds the BSE Hamiltonian as described in the text.